OpenStudy (anonymous):

the bill for your automobile repairs was $375.00 THe cost of labor is $50.00 per hou. the cost for materials is $150. how many hours did the mechanic work on ur car?

OpenStudy (anonymous):

The cost of the automobile repairs, in dollars, is given by the function\[f(x)=50x+150\]where x the number of hours the mechanic worked on your car. If the overall cost was 375$ then\[375=50x+150\]\[50x=375-150\]\[50x=225\]\[x=\frac{225}{50}\]\[x=4.5\] So, the mechanic worked 4.5 hours on your car.

OpenStudy (anonymous):

wouldnt it be 4.4?

OpenStudy (anonymous):

since u divide 220 by 50 and u get 4.4

OpenStudy (anonymous):

It is\[\frac{225}{50}=\frac{25\times9}{25\times2}=\frac{9}{2}=4.5\]